OpenStudy (anonymous):

Andrew can paint the neighbor's house 3 times as fast as Bailey. The year Andrew and Bailey worked together, it took them 4 days. How long would it take each to paint the house?

OpenStudy (anonymous):

Equation: rate + rate = together rate 1/3x + 1/x = 1/2 ---- Multiply thru by 6x to get: 2 + 6 = 3x x = 8/3 days (time for andrew to do the job alone). 3x = 8 days (time for bailey to do the job alone).

OpenStudy (anonymous):

Let bailey's time be 3x days/job ;then his rate is 1/3x job/day ---------------------------- Andrew's time is x days/job ; andrew's rate is 1d/x job/day -------------------------------------- the year they worked together it took them 2 days. Together time = 2 days/job ; together rate = 1/2 job/day

OpenStudy (anonymous):

So would that make Andrew able to complete the house in 2 days and Bailey in 8?

OpenStudy (mertsj):

a = days it takes Andrew to paint the house 3a = days it takes Bailey to paint the house 1/a = part Andrew does in one day 1/3a = part Bailey does in one day. 1/4 = part they do together in 1 day. \[\frac{1}{a}+\frac{1}{3a}=\frac{1}{4}\] \[12+4=3a\] \[a=\frac{16}{3}\]